OpenStudy (anonymous):

find an equation of the line that passes through the points (4,6) and (4,7)

OpenStudy (anonymous):

first find slope m=(7-6)/(4-4) =undefined its vertical line because its slope is undefined. equation of vertical line passing through point (a,b) is x=a here a=4 can you find it now ?

OpenStudy (anonymous):

The slope : (7-6)/(4-4) = undefined which means there's no slope and the line is vertical. The line passes through the points (4,6) and (4,7). the x coordinates of the both points = 4 Therefore, the equation is x=4

OpenStudy (anonymous):

yes, thank you